An investor placed a total of $16,000 in two simple interest accounts. One account earned a simple interest rate of 5.9%, and the other earned a simple interest rate of 3.6%. The two accounts earned a total of $668. How much was invested in each account?
Answer by checkley75(3666)

.059X+.036(16000-X)=668
.059X+576-.036X=668
.023X=668-576
.023X=92
X=92/.023
X=4000 AMOUNT INVESTED @ 5.9%
16000-4000=12000 INVESTED @ 3.6%.
PROOF
.059*4000+.036*12000=668
236+432=668
668=668